OpenStudy (anonymous):

solve 3(5t+4)<13t-10

OpenStudy (anonymous):

I got t<-11 is that right?

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

ok do I need to write it differently or just leave it like that

jimthompson5910 (jim_thompson5910):

you could leave it like that or write it in set builder notation or interval notation

jimthompson5910 (jim_thompson5910):

depends on how your book wants the answer

OpenStudy (anonymous):

I think I'm supposed to use interval notation

jimthompson5910 (jim_thompson5910):

ok what do you get when you convert over

OpenStudy (anonymous):

uummm you write equal to all real number less than -11?

jimthompson5910 (jim_thompson5910):

you want to describe the interval of all numbers less than -11

jimthompson5910 (jim_thompson5910):

so if you wanted to described the interval from 5 to 23 (not including either endpoint), then you would write: (5, 23)

OpenStudy (anonymous):

(0,-11) ?

jimthompson5910 (jim_thompson5910):

0 is larger than -11 so that makes no sense

OpenStudy (anonymous):

I dont know then because I know its less than -11 but I dont have a number that its greater than...so do I just leave it open ended?

jimthompson5910 (jim_thompson5910):

how would I describe the interval of numbers larger than 5

jimthompson5910 (jim_thompson5910):

you would do something like 5, 6, 7, 8, ... so (5, 7) is one interval...but leaves out numbers like 8, 9, 10, etc (5, 23) is another interval, but numbers like 30, 40, 50 are left out this interval is open ended because there are infinitely many numbers that are greater than 5 so you would write \[\large\left(5,\infty \right )\]

jimthompson5910 (jim_thompson5910):

does that example help?

OpenStudy (anonymous):

so \[(-11,-\infty)\]

jimthompson5910 (jim_thompson5910):

close, but the negative infinity is always on the left because it's always smaller

jimthompson5910 (jim_thompson5910):

in my example, infinity is on the right since it's always larger

OpenStudy (anonymous):

ok \[(-\infty,-11)\]

jimthompson5910 (jim_thompson5910):

good, remember you're reading this from left to right (smaller to bigger) so that's why (0, -11) doesn't make any sense...however (-11, 0) does make sense

OpenStudy (anonymous):

but that doesn't work in my case because t is less than -11

jimthompson5910 (jim_thompson5910):

don't worry about (0, -11) and (-11, 0), they are just examples of what is good and bad notation for interval notation

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

the answer is \[\large\left(-\infty,-11 \right )\] and you already got that

jimthompson5910 (jim_thompson5910):

np